OpenStudy (anonymous):

Find the value of a if the graph of the exponential function y = a · 2^x passes through the point (3, 128). Select one: a. 2 b. 4 c. 8 d. 16

zepdrix (zepdrix):

Hey there :) \[\Large\rm y=a\cdot 2^x\]So we pass through the point \(\Large\rm (x,y)=(3,128)\) Let's plug it in and solve for a,

zepdrix (zepdrix):

\[\Large\rm 128=a\cdot 2^3\]\[\Large\rm 128=a\cdot 8\]Can you finish it up from there? :)

OpenStudy (anonymous):

wouldnt it be 128=a*6 not 8

zepdrix (zepdrix):

Ah let's be careful with our powers! ^^\[\Large\rm 2^3=\quad 2\cdot2\cdot2=\quad 4\cdot2=\quad 8\]

OpenStudy (anonymous):

oh yes so it would be D 16

zepdrix (zepdrix):

yay good job!